what is fraction of initial kinetic energy remaining.

Grade:6

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

Please attach the image of the question

Regards
Vikas TU
3 years ago
m1u1 = (m1 + m2)v.
Ef = ½ (m1 + m2)v2,
Ei = ½ m1u12.
Fraction of energy lost = (Ei - Ef)/Ei
= 1 - m1/(m1 + m2) = m2/(m1 + m2).